### v2

Abstract
We report the discovery of an isolated millisecond pulsar M15O (J2129+1210O) from the globular cluster M15 (NGC 7078) with a period of 11.06686 ms and a dispersion measure of 67.44 cm-3 pc. Its spin period is so close to the 10th harmonic of the bright pulsar M15A (11.06647 ms) that it was missed in previous pulsar search. We suggest adding the spectrum in the pulsar candidate diagnostic plot to identify new signals near the harmonics. M15O has the first spin frequency derivative and the second spin frequency derivative,being 1.79191(5) Hz and 3.3133(6) Hz , respectively. Its projected distance from the optical center of M15 is the closest among all the pulsars in M15. The origin can be something from the center of the massive and core-collapsed globular cluster M15.
keywords:
methods: data analysis - globular clusters: individual (M15) - pulsars: individual (M15O or J2129-1210O)1 Introduction
Pulsars are rotating neutron stars with strong dipole magnetic fields, firstly discovered in 1967 through radio observations (Hewish et al. 1968). The discoveries of pulsars provide a natural laboratory for extreme physical environments, such as binary pulsars (e.g., PSR B1913+16;Hulse & Taylor 1975), massive neutron stars (e.g., PSR J16142230;Demorest et al. 2010), and millisecond pulsars (MSPs; e.g., PSR B1937+21 ;Backer et al. 1982). Pulsar searches face challenges from increasing radio frequency interference (RFI), dispersion smearing due to free electrons in the interstellar medium, and signal scattering caused by irregularities in interstellar plasma (Bégin 2006). If the pulsars are in binary systems, the orbital motion and rotational drift of pulsars will affect the times of arrival(ToAs) of the pulses to reach the Earth, significantly increasing the detection complexity of pulsar signals (e.g., Anderson et al. 1990).
The prevalent pulsar search methods mainly utilize the Fourier domain acceleration search (FFT; e.g., PRESTO111https://github.com/scottransom/presto software suite; Ransom, Eikenberry, & Middleditch 2002) and the Fast Folding Algorithm search (e.g., RIPTIDE-FFA222https://riptide-ffa.readthedocs.io/en/latest/index.html;Morello et al. 2020. Both of these methods necessitate period folding of time domain files and subsequent signal identification through folded profiles (Ransom, Eikenberry, & Middleditch 2002). Cross-checking with multiple observations helps to verify candidates, to either confirm or deny a candidate. In addition, the stack search, which involves incoherently stacking power spectra across multiple observations, has been proven effective in enhancing the detection sensitivity of faint pulsars using archival data (e.g., Anderson 1993; Pan et al. 2016; Cadelano et al. 2018).
Globular clusters (GCs) are characterized by the high stellar densities in their cores and frequent dynamical interactions, such as exchange interactions and tidal trapping (Harris 1996, 2010 edition). GCs have become prolific sites for the discovery of MSPs (e.g., Pallanca et al. 2014). Pulsars in GCs are most prominently concentrated in compact spatial distributions, such as those in Terzan 5 (e.g., Ransom et al. 2005;Padmanabh et al. 2024) and 47 Tucanae (NGC 104; e.g., Freire et al. 2017), well within the half-power beamwidth of one beam for single-dish telescopes at L-band (e.g., ). In this environment, detection and differentiation of pulsar signals become somehow difficult with multiple pulsars especially with very bright pulsars as their signals may interfere with each other. For GCs hosting many known pulsars with similar dispersion measure(DM) values, this increases the difficulty of distinguishing a potential pulsar signal. Sometimes, the timing analysis has to be used to confirm a new pulsar as their spin periods can be very close (e.g., M3D and M3E; Li et al. 2024).
To determine whether a detected signal in the same direction corresponds to a known pulsar, we can compare the pulsar’s spin period and DM values. If there are pulsars with strong fluxes, the candidate list may contain numerous instances of their harmonics, especially for pulsars with long spin periods. If the period of a new pulsar is similar to or overlaps with that of a known pulsar or its harmonics, and their DM values are also close (owing to the concentrated DM distribution within the cluster), conventional folding algorithms struggle to separate the signals due to phase - alignment ambiguities. The standard screening process may fail to effectively distinguish between them, potentially leading to missing the new pulsar.
The Five-hundred-meter Aperture Spherical Radio Telescope (FAST, Nan et al. 2011) operates with the most sensitivity among all the single-dish radio telescopes due to its illuminated aperture of 300 m. There are 45 GCs in the FAST sky, and pulsars have been discovered in 17 of them. Among them, M15 (NGC 7078) has 14 pulsars ranked second after NGC 6517. It is one of the oldest core-collapsed clusters with a core density of and with a distance from the Sun of 10.4 kpc (Harris 1996, 2010 edition). A simulation predicts that it has the highest number of predicted pulsars among all the GCs in FAST sky (Yin et al. 2024). Therefore, M15 has become a high-priority target for searching pulsars.
In this study, we report the discovery of a new pulsar (M15O, J2129-1210O) by employing FFT and the stack search method. It hides near the 10th harmonic of PSR J2129+1210A (M15A), identified by visual inspection at the spectra directly, and was finally verified by timing analysis using TEMPO333http://nanograv.github.io/tempo/ (Nice et al. 2015). The subsequent Section 2 is for the details of the data and data reduction procedures. Section 3 describes the discovery and timing results of M15O. Section 4 presents the discussion regarding this new pulsar, and the conclusions provided in Section 5.
2 Data and Data Reduction
This study is based on the 19 observations of M15 conducted by FAST between November 2019 and February 2024. The longest single observation lasted up to 4.5 hr. The total integration time was 44 hr. A detailed summary of these archival data is provided in Table 1. These data were recorded by the central beam of the FAST 19-beam receiver in the L-band with 4096 frequency channels (each with a bandwidth of 0.122 MHz) and it covers the frequency range from 1.05 to 1.45 GHz (Jiang et al. 2019).
With 3′ beam size, all known pulsars in M15 can be observed in at least some of these data. We reprocessed these data with FFT and the stack search methods, aiming at searching for new pulsars in M15. In the latter method, the stacked power spectra at different DMs are obtained by incoherently summing the power spectra of all observations.
| Observation Date | MJD | (s) | Stack | FFT | Timing | SNR | Flux Density () |
| 2019.11.09 | 58796 | 4800 | 8.17 | 2.19 | |||
| 2020.08.30 | 59091 | 3300 | |||||
| 2020.09.22 | 59114 | 9000 | 6.60 | 1.29 | |||
| 2020.09.25 | 59117 | 14400 | 12.18 | 1.88 | |||
| 2020.12.21 | 59204 | 16200 | 9.18 | 1.34 | |||
| 2021.03.09 | 59282 | 8700 | 8.56 | 1.70 | |||
| 2022.01.02 | 59581 | 1800 | 5.21 | 2.28 | |||
| 2022.09.04 | 59826 | 1800 | |||||
| 2022.10.15 | 59867 | 3000 | |||||
| 2022.11.19 | 59902 | 6600 | 11.89 | 2.72 | |||
| 2022.12.20 | 59933 | 10200 | 9.35 | 1.72 | |||
| 2023.01.20 | 59964 | 10200 | 7.83 | 1.44 | |||
| 2023.02.20 | 59995 | 6600 | 4.65 | 1.06 | |||
| 2023.12.06 | 60284 | 13560 | 6.44 | 1.03 | |||
| 2023.12.19 | 60297 | 15840 | 9.47 | 1.40 | |||
| 2024.01.22 | 60331 | 9060 | 5.25 | 1.02 | |||
| 2024.01.23 | 60332 | 10800 | 6.44 | 1.15 | |||
| 2024.01.30 | 60339 | 6320 | 6.91 | 1.61 | |||
| 2024.02.17 | 60357 | 6375 | 7.65 | 1.78 | |||
| Total | 19 | 5 | 16 |
2.1 Fourier Domain Acceleration Search
The RFI in the data was labeled by applying the rfifind routine of PRESTO. The DM values of the known pulsars in M15 are distributed in the range 65.5 to 67.8 cm-3 pc, with the mean value approximated as 67.0 cm-3 pc. The prepsubband routine was used to dedisperse the PSRFITS data over a larger DM trial range from 63 to 70 cm-3 pc, with step size of 0.05 cm-3 pc. In order to reduce the interference caused by low-frequency noise, the rednoise routine was applied. The realfft routine was utilized to transform the dedispersed time series to the frequency domain. The accelsearch routine was used to conduct acceleration searches in the spectra. The isolated pulsars dominate the pulsar population in M15 (13 out of 14). Thus, a low value of 20 was selected ( means the largest Doppler drift of the pulse frequency; Ransom, Eikenberry, & Middleditch 2002). Finally, an updated version of JinglePulsar444https://github.com/jinglepulsar/jinglesifting (Pan et al. 2021) was applied to sift and fold promising candidates for visual inspection.
2.2 The Stack Search
The stack search also firstly was performed with the routines rfifind, prepsubband, and realfft in PRESTO in the same steps as the above FFT. With these steps, a total of 141 frequency domain series files can be obtained from different DM values. From each file, the real and imaginary parts are added in an orthogonal manner to form the power values. Power spectra from multiple datasets with the same DM values are added to improve the signal-to-noise ratio (SNR) of possible periodic candidates. We added the power of the harmonics (up to 16) sequentially to the fundamental frequency in order to mitigate the energy dispersion in the frequency domain. After each harmonic addition, the power spectrum was retained for subsequent analysis. Finally, signals with power values exceeding ( is the standard deviation of the power value) in each stacked power spectrum are extracted as candidates. These candidates were all drawn on the power spectrum, and then screened one by one by visual inspection. Moreover the optimal DM trial and spin frequency of each candidate were further used to fold. More details about the stack search method will be described in Dai et al. (2025 in preparation).
3 Results
We discovered three pulsars in total, namely PSR J2129+1210M, N and O (M15M, N, and O) with the FFT and stack search schemes. As the M15M and N were firstly discovered by the stacking method, we will mention them in the next paper (Dai et al. 2025, in preparation). In this paper, we focus on the discovery of M15O and the reason for missing it in previous searches.
With the FFT method, we detected all know pulsars in M15. In addition, a large number of high-frequency harmonics appear in the candidates, being harmonics mainly from M15A and M15B. By comparing their period and DM, it is difficult to distinguish whether two signals with similar periods are harmonics from known pulsars or new pulsars. Multiple pulsar signals can only be distinguished by precise timing measurements, especially in dense clusters (e.g., Terzan 5; Ransom et al. 2005). Eventually, a pulsar named M15O (J2129-1210O) with a period of 11.06687 ms was detected. M15O stands out with a relatively high SNR of 11.0. Its spin period is very close to the 10th harmonic of M15A (11.06647ms), with a difference of only 0.004 ms. With the 10th harmonic of M15A, the power spectrum of this pulsar is displayed in Figure 1.
The folded plot of M15O using the PRESTO prepfold routine is shown in Figure 2. The anomalous morphology in the time domain depicted in Figure 2 (left panel) arises from phase interference between M15O and harmonics of M15A. The diagonal striations in the time series are from the M15A’s harmonics. To further validate our findings, we chopped the M15A’s signal in the frequency domain using PRESTO’s simple_zapbirds.py. Then, a time domain series was generated by the inverse Fourier transformation using realfft. After folding that time series, the M15O signal remains, yet without the harmonic of M15A.
| Pulsar | J2129-1210O |
| Right Ascension, (J2000) | 21:29:58.2999(9) |
| Declination, (J2000) | +12:10:00.97(2) |
| Spin Frequency, (s-1) | 90.3597874705(1) |
| First Spin Frequency derivative, (Hz s-1) | 1.79191(5) |
| Second Spin Frequency derivative, (Hz s-2) | 3.3133(6) |
| Reference Epoch (MJD) | 60297.270620 |
| Start of Timing Data (MJD) | 58404.483 |
| End of Timing Data (MJD) | 60357.276 |
| Dispersion Measure, DM (pc cm-3) | 67.44(1) |
| Number of TOAs | 65 |
| Residuals RMS (s) | 108.00 |
The timing analysis was performed with the TEMPO package. We employed the Dracula555https://github.com/pfreire163/Dracula algorithm (Freire & Ridolfi 2018) to determine the precise rotation counts of pulsars. The solution (see Table 2) returned from Dracula was consistent with the results obtained by manually removing the arbitrary phase between epochs. The pulse profile and timing residuals of M15O are presented in Figure 3. The position of M15O has a projected distance of 0.37′′ from the cluster center and 0.81′′ from M15A (Figure 4), making it the closest known pulsar to the M15 core in the projected sky. This confirms that M15O and M15A are spatially distinct.
4 Discussion
4.1 Possible Reasons of Missing M15O in Previous Data Reduction
We suggest that the key point in M15O identification is from the spectra. As the peak near the 10th harmonic of M15A was seen, it is clear that such a signal can be something different from the harmonic itself. The M15O is somehow easy to be detected by PRESTO based on typical pulsar search pipeline (Yu et al. 2020). We looked back to the previous pulsar search results and found that M15O should be discovered in January of 2021 or even earlier. There are two possible reasons as bellows.
Firstly, periodic proximity introduces selection biases in candidate evaluation. The harmonics of M15A can be the main affection in the identification of M15O. The discovery plot (Figure 2, left) is not so useful in this situation that the M15A signal dominates the features in all the panels. A quick test showed that at least 50% of experienced persons (colleagues around us, including experienced persons who found numbers of pulsars) were not agree that a new pulsar signal can be see in such a plot. Crucially, definitive confirmation of M15O required phase coherent timing analysis (position differences) rather than conventional a re-detection, indicating that this signal is quite easy to be mixed with the harmonics of M15A.
Secondly, M15O’s intrinsic faintness presents detection challenges even in FAST data. Its SNR is around 11 in the power spectrum, similar to the pulsar M15M and M15N which were discovered by stacking 44 hours FAST observational data. Thanks to M15O’s relatively narrow pulse profile, making it easily to be distinguished from the noise level.
The radiometer equation is widely used to calculate the minimum detectable flux density of a pulsar (Lorimer & Kramer 2012), as expressed by:
where represents the sampling efficiency of the 8-bit recording system. The system temperature is 20 K, the antenna gain is 16 K Jy-1, and the number of polarizations is 2. The bandwidth is set to 400 MHz (Jiang et al. 2020). The is 0.15 for M15O. The estimated flux densities of M15O range from 1.02 to 2.72 , as shown in Table 1.
4.2 The Large F1 and Measured F2 of M15O
A phase connected timing solution of M15O could not be achieved by fitting only the spin frequency (F0) and first spin frequency derivative (F1). However, the inclusion of the second spin frequency derivative (F2) enabled us to obtain a single solution that connects all the available data (see Table 2). A similar situation is observed for M15I. It has a large F1 and also requires the inclusion of F2, but no single timing solution has been obtained so far. Based on the timing results of M15O, the F1 is remarkably large, second only to that of M15D among all known pulsars in M15 (see Table 3). Moreover, M15O is the only pulsar in this cluster for which the F2 has been successfully measured. M15O is located extremely close to the cluster center, with a projected distance of . These variations are likely caused by the complex gravitational potential near the center of M15, which may be due to the high-density interstellar environment or the orbital motion around a massive object. If the latter is the case, we may estimate the orbital period to be on the order of decades based on F1 and F2. Our observations with FAST have already spanned six years, and we plan to continue monitoring M15 with FAST. It is possible that future observations may reveal the nature of the very large F2 of either M15O or M15I.
| Pulsar | Spin Frequency (s-1) | F1 (Hz s-1) | F2 (Hz s-2) |
| J2129+1210O | 90.3597874705(1) | 3.3133(6) | |
| J2129+1210A | 9.0363060726718(5) | …… | |
| J2129+1210B | 17.814819156216(5) | …… | |
| J2129+1210C | 32.7554227003(4) | …… | |
| J2129+1210D | 208.21173107090(3) | …… | |
| J2129+1210E | 214.987399714593(8) | …… | |
| J2129+1210F | 248.32119058101(2) | …… | |
| J2129+1210G | 26.553254549(6) | …… | |
| J2129+1210H | 148.2932725188(7) | …… | |
| J2129+1210I* | 195.228710(2) | …… | …… |
| J2129+1210J | 84.44174596102(2) | …… | |
| J2129+1210K | 0.51855097775(1) | …… | |
| J2129+1210L | 0.25247957713(1) | …… | |
| J2129+1210M | 206.77469911787(9) | …… | |
| J2129+1210N* | 107.64786(1) | …… | …… |
4.3 The Possibility of New Pulsars Hidden in the Known Pulsar Signals
Pulsars observed in the same GC generally are located near the center and exhibit similar DM values. We conducted a statistical analysis of pulsars ( ones) outside of GCs from the PSRCAT666https://www.atnf.csiro.au/research/pulsar/psrcat/, to see if any pair of pulsars can be included in a 5′ diameter circle and the DM difference between them is less than 5 cm-3 pc. The results indicate that no such combinations were found. So, it seems that currently such hidden pulsar may not be missed in the Galactic field pulsar surveys. On the other hand, it is worth noting that due to FAST’s sensitivity, even more hidden pulsars may occur in GC data, even though M15O is the only case until now.
As radio telescope sensitivity has gradually increased in the past ten years, the number of detected pulsars keeps rising. The pulsars discovered by FAST exceeded 1000 at the end of 2024, futher increasing the pulsar density in the Galactic filed. The probability of encountering such cases, whether in GCs or on the Galactic disk, is also increasing. We suggest that the spectra obtained from the FFT should be presented in a diagnostic plot. This will assist the identification of pulsar harmonics or new potential signals. We are looking forward to seeing if more can be identified through spectra-stacking searches on other GC data from FAST. In GC pulsars, we are checking the possible new pulsars hidden in the known pulsars harmonics, and will report the results in an upcoming paper focusing on the results from other GCs.
5 Conclusions
The conclusions are as follows:
(1) An isolated MSP, namely M15O (J2129-1210O), with a period of 11.06687 ms and a DM of 67.44 cm-3 pc was discovered. Its spin period is very close to the 10th harmonic of the bright and previously known pulsar M15A (11.06647 ms).
(2) The discovery of M15O is due to the stack search method. To obtain pulsar candidates, each peak in the stacked spectrum was checked carefully. The timing analysis was used to confirm this new pulsar. This discovery validates the effectiveness of the stack search and improved candidate screening strategies in mitigating harmonic interference.
(3) We obtained the phase connected timing solution for M15O. It has large F1 and measurable F2, being Hz s-1 and Hz s-2, respectively. Its projected distance from the optical center of M15 is the closest among all the known pulsars in M15.
(4) We suggest adding the power spectrum in the pulsar candidate diagnostic plot to identify possible new signals near the harmonics of previously known pulsars.
Acknowledgements.
We thank the referee’s valuable suggestions, which helped to refine our paper. We thank Paulo C. C. Freire and Scott M. Ransom for valuable discussions and suggestions during the validation of M15O, and Paulo C. C. Freire for helping us obtain the ephemeris. This work is supported by the National Key R&D Program of China No. 2022YFC2205202, No. 2020SKA0120100 and the National Natural Science Foundation of China (NSFC, Grant Nos. 12373032, 12003047, 11773041, U2031119, 12173052, and 12173053. Both Lei Qian and Zhichen Pan were supported by the Youth Innovation Promotion Association of CAS (id. 2018075, Y2022027 and 2023064), and the CAS ”Light of West China” Program. Liyun Zhang has been supported by the Science and Technology Program of Guizhou Province under project No.QKHPTRC-ZDSYS[2023]003 and QKHFQ[2023]003. This work made use of the data from FAST (Five-hundred-meter Aperture Spherical radio Telescope) (https://cstr.cn/31116.02.FAST). FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences.References
- Anderson et al. (1990) Anderson S. B., Gorham P. W., Kulkarni S. R., Prince T. A., Wolszczan A., 1990, Natur, 346, 42. doi:10.1038/346042a0
- Anderson (1993) Anderson S. B., 1993, PhDT
- Backer et al. (1982) Backer D. C., Kulkarni S. R., Heiles C., Davis M. M., Goss W. M., 1982, Natur, 300, 615. doi:10.1038/300615a0
- Bégin (2006) Bégin S., 2006, MsT, 5
- Cadelano et al. (2018) Cadelano M., Ransom S. M., Freire P. C. C., Ferraro F. R., Hessels J. W. T., Lanzoni B., Pallanca C., et al., 2018, ApJ, 855, 125. doi:10.3847/1538-4357/aaac2a
- Demorest et al. (2010) Demorest P. B., Pennucci T., Ransom S. M., Roberts M. S. E., Hessels J. W. T., 2010, Natur, 467, 1081. doi:10.1038/nature09466
- Freire et al. (2017) Freire P. C. C., Ridolfi A., Kramer M., Jordan C., Manchester R. N., Torne P., Sarkissian J., et al., 2017, MNRAS, 471, 857. doi:10.1093/mnras/stx1533
- Freire & Ridolfi (2018) Freire P. C. C., Ridolfi A., 2018, MNRAS, 476, 4794. doi:10.1093/mnras/sty524
- Hewish et al. (1968) Hewish A., Bell S. J., Pilkington J. D. H., Scott P. F., Collins R. A., 1968, Natur, 217, 709. doi:10.1038/217709a0
- Heinke, Cohn, & Lugger (2009) Heinke C. O., Cohn H. N., Lugger P. M., 2009, ApJ, 692, 584. doi:10.1088/0004-637X/692/1/584
- Hulse & Taylor (1975) Hulse R. A., Taylor J. H., 1975, ApJL, 195, L51. doi:10.1086/181708
- Harris (1996) Harris W. E., 1996, AJ, 112, 1487. doi:10.1086/118116
- Jiang et al. (2020) Jiang P., Tang N.-Y., Hou L.-G., Liu M.-T., Krčo M., Qian L., Sun J.-H., et al., 2020, RAA, 20, 064. doi:10.1088/1674-4527/20/5/64
- Jiang et al. (2019) Jiang P., Yue Y., Gan H., Yao R., Li H., Pan G., Sun J., et al., 2019, SCPMA, 62, 959502. doi:10.1007/s11433-018-9376-1
- Li et al. (2024) Li B., Zhang L.-. yun ., Yao J., Yin D., Eatough R. P., Li M., Li Y., et al., 2024, ApJ, 972, 43. doi:10.3847/1538-4357/ad5a82
- Lorimer & Kramer (2012) Lorimer D. R., Kramer M., 2012, hpa..book
- Morello et al. (2020) Morello V., Barr E. D., Stappers B. W., Keane E. F., Lyne A. G., 2020, MNRAS, 497, 4654. doi:10.1093/mnras/staa2291
- Nice et al. (2015) Nice D., Demorest P., Stairs I., Manchester R., Taylor J., Peters W., Weisberg J., et al., 2015, ascl.soft. ascl:1509.002
- Nan et al. (2011) Nan R., Li D., Jin C., Wang Q., Zhu L., Zhu W., Zhang H., et al., 2011, IJMPD, 20, 989. doi:10.1142/S0218271811019335
- Pallanca et al. (2014) Pallanca C., Ransom S. M., Ferraro F. R., Dalessandro E., Lanzoni B., Hessels J. W. T., Stairs I., et al., 2014, ApJ, 795, 29. doi:10.1088/0004-637X/795/1/29
- Padmanabh et al. (2024) Padmanabh P. V., Ransom S. M., Freire P. C. C., Ridolfi A., Taylor J. D., Choza C., Clark C. J., et al., 2024, A&A, 686, A166. doi:10.1051/0004-6361/202449303
- Pan et al. (2016) Pan Z., Hobbs G., Li D., Ridolfi A., Wang P., Freire P., 2016, MNRAS, 459, L26. doi:10.1093/mnrasl/slw037
- Pan et al. (2021) Pan Z., Ma X.-Y., Qian L., Wang L., Yan Z., Luo J.-T., Ransom S. M., et al., 2021, RAA, 21, 143. doi:10.1088/1674-4527/21/6/143
- Pan et al. (2021) Pan Z., Qian L., Ma X., Liu K., Wang L., Luo J., Yan Z., et al., 2021, ApJL, 915, L28. doi:10.3847/2041-8213/ac0bbd
- Wu et al. (2024) Wu Y., Pan Z., Qian L., Ransom S. M., Eatough R. P., Wang B., Freire P. C. C., et al., 2024, ApJL, 974, L23. doi:10.3847/2041-8213/ad7b9e
- Ransom, Eikenberry, & Middleditch (2002) Ransom S. M., Eikenberry S. S., Middleditch J., 2002, AJ, 124, 1788. doi:10.1086/342285
- Ransom et al. (2005) Ransom S. M., Hessels J. W. T., Stairs I. H., Freire P. C. C., Camilo F., Kaspi V. M., Kaplan D. L., 2005, Sci, 307, 892. doi:10.1126/science.1108632
- Yin et al. (2024) Yin D., Zhang L.-. yun ., Qian L., Eatough R. P., Li B., Lorimer D. R., Dai Y., et al., 2024, ApJL, 969, L7. doi:10.3847/2041-8213/ad534e
- Yu et al. (2020) Yu Q.-Y., Pan Z.-C., Qian L., Wang S., Yue Y.-L., Huang M.-L., Hao Q.-L., et al., 2020, RAA, 20, 091. doi:10.1088/1674-4527/20/6/91